### v2

# ReLI: A Language-Agnostic Approach to Human-Robot Interaction

Abstract
Adapting autonomous agents to industrial, domestic, and other daily tasks is currently gaining momentum. However, in the global or cross-lingual application contexts, ensuring effective interaction with the environment and executing unrestricted human task-specified instructions in diverse languages remains an unsolved problem. To address this challenge, we propose ReLI, a language-agnostic framework designed to enable autonomous agents to converse naturally, semantically reason about the environment, and to perform downstream tasks, regardless of the task instruction’s linguistic origin. First, we ground large-scale pre-trained foundation models and transform them into language-to-action models that can directly provide common-sense reasoning and high-level robot control through natural, free-flow human-robot conversational interactions. Further, we perform cross-lingual grounding of the models to ensure that ReLI generalises across the global languages. To demonstrate the ReLI’s robustness, we conducted extensive simulated and real-world experiments on various short- and long-horizon tasks, including zero-shot and few-shot spatial navigation, scene information retrieval, and query-oriented tasks. We benchmarked the performance on 140 languages involving over 70K multi-turn conversations. On average, ReLI achieved over 90%$\pm$0.2 accuracy in cross-lingual instruction parsing and task execution success rates. These results demonstrate the ReLI’s potential to enhance natural human-robot interaction in the real world while championing linguistic diversity. Demonstrations and resources will be publicly available at https://linusnep.github.io/ReLI/.
I Introduction
Nowadays, autonomous agents are increasingly being employed for various real-world tasks, including industrial inspection, domestic chores, and other daily tasks. However, as the challenges presented to these agents become more intricate and the environments they operate in grow more unpredictable [6, 60, 77] and linguistically diverse, there arises a clear need for more effective and language-agnostic human-robot interaction (HRI) mechanisms [78, 84]. Until now, language has posed a formidable obstacle to achieving truly universal and realistic natural human-robot collaboration in real-world environments [6, 26, 61]. Most autonomous agents have been constrained by unilateral linguistically specific training, often restricted to widely spoken (high-resource) languages such as Chinese, French, and English. Therefore, to preserve linguistic diversity and promote inclusive and accessible HRI [5, 58, 87] in the real world, enabling autonomous agents to converse across multiple languages is essential.
Figure 1: An illustration of how ReLI enables autonomous agents to perform both short- and long-horizon tasks. (a) A natural language instruction $c\in\mathcal{C}_{T}$ is given regardless of the language $\ell\in\mathcal{L}$ of the task instruction. In (b) and (c), ReLI reasons over the task instruction and autoregressively generates a sequence of action plans, i.e., $Action_{1},Action_{2},\dots,Action_{7}$ that accomplishes the given task. (d) It then seeks the user’s consent for these action plans (i.e., in the case of multistep actionable commands) before transmitting them to the robot’s controller for physical execution. (e) If the user affirms, the parsed instructions will be executed; otherwise, they will be discarded. We refer the reader to Section III for the formal details.
The HRI community has been instrumental in proffering solutions to these long-standing goals. However, despite the remarkable progress from the community, a significant proportion of the existing HRI frameworks [43, 35, 1] and benchmarks [65, 64, 4, 83] predominantly cater for high-resource languages [80, 21, 68, 42, 31] or often rely on the complex robot teleoperation interfaces [96, 50]. To our knowledge, there exists no framework that enables autonomous agents to converse naturally, interact with their environment, and perform downstream tasks regardless of the language of task instructions. Consequently, these linguistic and technical barriers, imposed by the reliance on unilateral language paradigms, can disproportionately impact the usability and accessibility of natural language-conditioned robotic systems.
Prompted by these challenges, we propose, Regardless of the Language of task Instructions (ReLI). ReLI is a free-form multilanguage-to-action framework, designed to push the boundaries of natural human-robot interaction in the real world. In particular, we introduced a language-agnostic approach that accommodates diverse linguistic backgrounds, including endangered languages, Creoles and Vernaculars, e.g., African Pidgin, USA Cherokee, etc., and various levels of technical expertise in HRI. To achieve these novel objectives, we extensively exploit the inherent cross-lingual generalisation capabilities [107, 95] of large-scale pre-trained foundation models, e.g., GPT-4o [66], to capture semantic and syntactic aspects across languages without explicit supervision for each language, data collection, and model retraining.
Fig. 1, illustrates how ReLI empowers autonomous agents to solve both short- and long-horizon tasks simply from human-specified natural language commands. Overall, ReLI capabilities are broad and include, but are not limited to, the ability to enable autonomous agents to (i) perform language-conditioned tasks over both short and long horizons, and (ii) execute the task instructions regardless of their linguistic origin. Thus, this work makes the following key contributions:
  • We introduce ReLI, a robust language-agnostic and inclusive framework for real-world human-robot collaboration tasks. Unlike the existing approaches that depend on code-level methods [48] or unilingual high-resource languages [8, 61, 1, 24, 79], ReLI is the first language-conditioned HRI framework to abstract natural free-form human instructions into robot actionable commands, regardless of the language of the task instruction.
  • We conducted extensive real-world and simulated experiments with ReLI on several short and long-horizon tasks, including zero and few-shot spatial navigation, scene information retrieval, and query-oriented tasks.
  • We benchmarked ReLI’s multilingual performance on $140$ human-spoken languages, involving over $70$K multi-turn conversations. In all the tested languages, ReLI achieved on average $90\%\pm 0.2$ accuracy in multilingual instruction parsing and task execution success rate.
  • ReLI generalises across different command input modalities and scenarios to allow off-the-shelf human-robot interaction regardless of technical expertise.
II Related Works
The last few years have witnessed tremendous advancement in generative AI [59, 56, 52] and natural language processing (NLP) [11, 33, 30, 22]. This surge, primarily driven by large language models (LLMs) [66, 13, 88, 93] has revolutionised the way intelligent systems process and interpret human instructions [20, 99, 75, 104]. LLMs, trained on extensive corpora sourced from the web [69], are typically autoregressive transformer-based architectures [94, 19]. In principle, given an input sequence $c=(c_{1},c_{2},\dots,c_{T})\in\mathcal{C}_{T}$, where $\mathcal{C}_{T}$ represents the space of all possible user commands, these models predict the corresponding output tokens $y=(y_{1},y_{2},\dots,y_{T})\in\mathcal{Y}_{T}$ with $\mathcal{Y}_{T}$ being the space of all possible outputs sequences of sequence length $T$. They employ the chain rule of probability to factorise the joint distribution over the output sequence, as illustrated in Eq. (1), ensuring context-sensitive decoding at each step, where $\theta$ represents the learned model parameters:
$\begin{split}p_{\theta}(y_{1},y_{2},\dots,y_{T}\mid c)&=p_{\theta}(y_{1}\mid c)\cdot p_{\theta}(y_{2}\mid y_{1},\;c)\dots\,,\\ p_{\theta}(y_{T}\mid y_{1:T-1},\;c)&=\prod_{t=1}^{T}p_{\theta}(y_{t}\mid y_{1:t-1},\;c).\end{split}$ (1)
Although these LLMs were originally designed as powerful language processing engines [109, 9], their quantitative and qualitative abilities [86], including multilingual capabilities, have been rigorously evaluated by independent third parties. In recent years, several works [21, 81, 2, 44, 106] have shown that these models can achieve exceptional generalisation across languages, beyond the high-resource languages that traditionally dominate the natural language processing benchmarks [16, 49, 89]. Thus, this multilingual prowess makes these models compelling candidates for communication and instruction in linguistically heterogeneous environments.
On the other hand, visual language models (VLMs) [73, 46] pre-trained on large-scale image-text pairs have emerged as a groundbreaking approach to integrate visual and textual modalities. These models leverage the synergies between visual data and natural language to enable agents to semantically and effectively reason about their task environment, where the traditional computer vision models fumble. In principle, they employ contrastive learning techniques [36] to align visual features with the corresponding textual descriptions.
In the field of robotics, the integration of VLMs with LLMs has unlocked several avenues for multimodal reasoning [98, 25, 40, 102] and task grounding [1, 62, 92]. Translating from language to real-world action is the most common form of grounding robotic affordances in recent years [51, 100, 53]. Recent works [32, 82, 61, 103, 8] have demonstrated that with VLMs and LLMs combined, robots can perceive and perform long-horizon tasks specified in free-form natural language in a manner akin to human cognition.
However, despite these advances, grounding these models to multilingual robotic affordances remains an open challenge. To date, most language-instructable [1, 3, 61] and visual-language-conditioned [17, 37, 47, 101] robotic frameworks have primarily focused on grounding unilingual instructions or limited high-resource languages [41] into low-level robotic actions. These approaches often struggle with the complexities of cross-lingual instructions and intricate task specifications, as they are not designed to translate natural language commands from diverse linguistic backgrounds into robotic actions.
Consequently, while these approaches have achieved impressive results in grounding foundation models in real-world robotic affordances, their inability to handle diverse multilingual instructions limits their adaptability in cross-linguistic operational domains. In this work, we tackled these challenges. In particular, we proposed a novel foundation model-driven language-agnostic HRI approach that combines the inherent strengths of language and visual foundation models. With the combined strengths, we realised a new universal approach to human-robot interaction, one where, regardless of the conversation modality or the language of the task instruction, the conversation is the robot’s command.
III Method
III-A Problem Description
We address the problem of grounding multilingual natural language instructions, expressed in free-form human languages, into robotic affordances. Specifically, we consider (i) high-level user-instructible linguistic commands $\mathcal{C}_{T}$ in a language $\ell\in\mathcal{L}$ generalisable by the state-of-the-art LLMs (e.g., GPT-4o [66], DeepSeek [13], etc.), and (ii) a set of high-dimensional observations $\mathcal{V}_{s}$ (e.g., synchronised RGB-D data, odometry) from the robot’s onboard perception sensors, that captures the state of the environment. Our primary objective is to learn the mapping function $\mathcal{F}_{LLM}:\mathcal{C}_{T}\times\mathcal{V}_{s}\mapsto\mathcal{A}$ that transforms the instruction and the current visual context into a sequence of physically robot executable actions $\mathcal{A}$. Critically, we require the resulting output $\mathcal{F}_{LLM}(.)$ to generalise across languages, $\mathcal{L}$, to allow task instructions to be interpreted and executed regardless of their linguistic origin.
Figure 2: Overview of ReLI’s architecture. For user’s commands in languages generalisable by the state-of-the-art LLMs, we decompose ReLI functionality into four main components that involve: (a) language detection and transcription, (b) instruction reasoning, processing and instruction-to-action parsing, (c) knowledge-based visuo-lingual and spatial grounding, and (d) real-world robot control and action execution. Refer to Section III for the details.
In order to accomplish these high-level objectives, we decomposed our approach into four architectural taxonomies based on individual functions. Fig. 2 illustrates our framework’s overview and the architectural decomposition. First, we present the user access point, where the input modalities and task instructions are detected and transcribed (i.e., in the case of vocal or audio instructions $c^{v}$) into textual representations (Subsection III-B). Second, we exploit the capabilities of a large-scale pre-trained LLM to reason over the high-level natural language instructions and abstract them into robot-actionable commands (Subsection III-C). Third, we provide a visual and semantic understanding of the robot’s task environment through a contrastive language image pre-training model alongside a supervised computer vision model (Subsection III-D). Finally, we abstract the high-level understanding from our action decision pipeline (Subsection III-C) into the actual physical robot actions (Subsection III-E).
III-B Bidirectional Interaction Interface
The bidirectional interaction interface (I2Face), shown at the top-left of Fig. 2 serves as the user’s primary access point to our framework. We developed the interface utilizing the Tkinter libraries [54] and integrated it into our framework through the robot operating system (ROS) [72] message-passing communication protocol [^1] Specifically, we employed the standard ROS [72] publish & subscribe communication mechanism, in which the I2Face and the DnPP exchange messages in bidirectional approach. User inputs (including the transcribed textual representation, $\hat{c}^{t}\in\mathcal{C}_{\text{audio}}$) are published to the DnPP, and the responses are subsequently subscribed to and displayed by I2Face. In our design, we employed this event-driven approach to ensure that user actions, such as sending a message or issuing a command, trigger corresponding interface updates or direct command publications to the DnPP..
User-provided natural language instructions can arise through two primary input modalities, namely plain text $c^{t}\in\mathcal{C}_{\text{text}}$ and audio or vocal instructions $c^{v}\in\mathcal{C}_{\text{audio}}$. To accommodate both modalities, we developed a method that consolidates the user’s instructions such that all commands converge into a single text-based domain suitable for further linguistic processing. For users without direct access to the I2Face (e.g., for inputting textual instructions), we introduced an automatic speech recognition (ASR) method [10, 74] that captures high-level auditory input and transcribes it into textual representations. We express this transformation as $\hat{c}^{t}=\mathrm{ASR}\bigl{(}c^{v},\ell_{i}\bigr{)}$, where $\ell_{i}$ denotes a finite set $\{\ell_{1},\ell_{2},\dots,\ell_{n}\}$ of LLM-generalisable languages.
Moreover, our framework also provides the users with the capability to manually select the language $\ell_{i}$ to configure the ASR models and dictionaries such that $\hat{c}^{t}$ accurately reflects the source language’s acoustic and syntactic properties. With the instruction transcribed into textual representation, we map them to the decision and parsing pipeline (DnPP), Section III-C, where interpretation and action derivation occur. Fig. 3 illustrates the I2Face, showing how ReLI can dynamically adapt to any language of task instruction.
III-C Command Decision and Parsing
Fig. 2 (middle left) illustrates the action decision and parsing pipeline (DnPP). We conceptualize our approach for multilingual language-to-action grounding as a decision process wherein an arbitrary linguistic command $c\in\mathcal{C}_{T}$, specified in language $\ell\in\mathcal{L}$, is transformed into a sequence of robot-executable commands. For these transformation, we extensively exploit the chain-of-thought reasoning [81, 108, 97] of pre-trained LLMs to map $c$ into an equivalent sequence of robot-executable high-level instructions, $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{k}\}$. Each $a_{i}$ corresponds to an atomic sub-instruction derived from the LLM’s interpretation of $c$. Specifically, we modelled the action decision process as an LLM mapping task $\mathcal{F}_{\mathrm{LLM}}$ that, given $c\in\mathcal{C}_{T}$, infers a high-level interpretation $r\in\mathcal{R}_{\text{int}}=\mathcal{F}_{\mathrm{LLM}}(c)$ of the user’s intent. For a set of LLM-generalisable languages, and user-provided commands in the language $\ell$, we define a latent variable model that assigns a probability distribution over an action sequence $\mathcal{A}$ as depicted in Eq. (2). The distribution is marginalized over $r\in\mathcal{R}_{\text{int}}$, where $\theta$ denotes the LLM parameters:
$p_{\theta}(\mathcal{A}\mid c,\,\ell)=\sum_{r\in\mathcal{R}_{\text{int}}}p_{\theta}(\mathcal{A}\mid c,\,\ell,\,r)\,p_{\theta}(r\mid c,\,\ell).$ (2)
We further factorize the conditional term in Eq. (2) autoregressively (see Eq. (3)) to ensure that each action token $a_{i}$ is generated in context, conditioned on prior actions $\{a_{1},\dots,a_{i-1}\}$, the command $c$, the language $\ell$, and the LLM’s high-level interpretation $r$ of the intended tasks:
$p_{\theta}(\mathcal{A}\mid c,\,\ell,\,r)=\prod_{i=1}^{k}p_{\theta}\bigl{(}a_{i}\mid a_{<i},\,c,\,\ell,\,r).$ (3)
To produce a deterministically structured action plan, we employ a rule-based commands parser $\mathcal{P}$ to translate $r$ into a set of low-level actionable primitives, as depicted in Eq. (4):
$\mathcal{P}(r)=\{\mathcal{A}_{1}(\phi_{1}),\dots,\mathcal{A}_{n}(\phi_{n})\}=\{\mathcal{A}_{j}(\phi_{j})\}_{j=1}^{n},\,$ (4)
where each discrete action token $\mathcal{A}_{j}$ is drawn from predefined action vocabulary (e.g., forward, turn right, pause, etc.) with $n\geq k$ to account for potential high-level actions that may require expansion to multiple primitives (e.g., “move in a square pattern”), and $\phi_{j}\in\mathbb{R}^{{m}_{j}}$ parameterizes the action primitives (e.g., distance ($m$), angle (), speed ($m/s$), etc.).
To handle multilingual inputs, we leveraged the LLM’s language-agnostic and multilingual capabilities to ensure ReLI’s generalisation to diverse languages. Concretely, when a user provides an instruction, first, we define a language detection $\mathcal{L}_{\text{dect}}$ pipeline, which infers the language $\ell$ of the given instruction, i.e., $\ell=\mathcal{L}_{\text{dect}}(c)$. However, if the user explicitly sets $\ell$ via the canonical language codes pre-defined at the I2Face (Subsection III-B), $\mathcal{L}_{\text{dect}}$ is automatically bypassed, and the internal language is configured according to the chosen $\ell$. Once $\ell$ is established, we then condition the output distributions of the LLM’s description from Eq. (3) such that the instruction in $\ell$ is appropriately parsed and interpreted according to the language’s syntactic and lexical norms. Simultaneously, we update the internal user-language state to the current $\ell$ (see Fig. 3). This practically keeps the conversation coherent and ensures that any subsequent actions are dynamically updated in the same language.
Figure 3: ReLI employs a dynamic and event-driven architecture where each user’s language input triggers a corresponding response. Additionally, action execution updates are provided in the same language as the input to ensure seamless bidirectional communication and interaction.
Additionally, to ensure that the interpreted action plans accurately reflects the user’s intent, especially for complex long-horizon tasks, we introduce a security confirmation mechanism that prompts the user to approve or reject the generated action plans before being deployed for execution. We modelled the confirmation as a binary decision problem, $\rho_{d}\in\{0,1\}$, inferred by applying a learned linear classifier $\mathbf{v}$ to an embedding $\psi(r)$ of the LLM’s interpretation $r$, as shown in Eq. (5):
$\rho_{d}=\begin{cases}1&\text{if }\mathbf{v}^{\top}\psi(r)>0\implies\text{execute the plan}\\ 0&\text{otherwise}\implies\text{discard the plan}\end{cases}.$ (5)
We classified the user responses into positive and negative templates, instead of simply yes or no. For example, positive responses include phrases like “sure, go ahead”, while negative responses might be “not correct, cancel”, etc.
III-D Visuo-lingual Perception and Object Localization
The ReLI’s visuo-lingual pipeline (bottom of Fig. 2) relies on foundation vision-language models, e.g., CLIP [73] and zero-shot segmentation models (SAM [38]), augmented with geometric depth fusion and uncertainty-aware classification to ground linguistic references into spatially localised entities in the robot’s operational environment. Formally, let $\mathcal{V}_{s}=\{(\mathcal{I}_{t},\mathcal{D}_{t},u_{t})\}_{t=1}^{T}$ be the sequence of time-synchronized RGB-D frames and odometry signals from the robot’s observation sensors, where $\mathcal{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ is the stream of RGB frames, $\mathcal{D}_{t}\in\mathbb{R}^{H\times W}$ is the corresponding depth map, and $u_{t}$ encodes the transformations in the robot’s local frame at time $t$. For each $\mathcal{I}_{t}$, we employ SAM [38] to generate $N$ candidate masks $\{\mathcal{M}_{i}\}_{i=1}^{N}$ through prompt-guided and automatic segmentation.
For each mask $\mathcal{M}_{i}$, we employ convex hull analysis to evaluate the quality. The ratio of the mask area to the convex hull area $\mathbf{Q}(\mathcal{M}_{i})$ determines its validity, and low-quality masks with $\mathbf{Q}(\mathcal{M}_{i})<\tau_{q}$ are discarded, where $\tau_{q}$ is the quality threshold. For retained valid masks with $\mathbf{Q}(\mathcal{M}_{i})>\tau_{q}$, we isolate the image regions and encode them into CLIP’s joint visual-textual space. We then compare the visual embeddings $\mathcal{V}_{i}=f_{\text{visual}}(\mathcal{I}_{t}\odot\mathcal{M}_{i})$ to text embeddings $t_{j}=f_{\text{text}}(d_{j})$ of candidate labels $\{d_{j}\}_{j=1}^{M}$ via $S_{ij}=\cos(\mathcal{V}_{i},t_{j}),$ being the similarity score. Further, we apply a softmax function with learned temperature parameter $\mathbf{T}$ to yield a probability distribution over classes, as:
$p(d_{j}\mid\mathcal{M}_{i})=\frac{\exp\!\bigl{(}\tau\,S_{ij}\bigr{)}}{\sum_{k=1}^{M}\exp\!\bigl{(}\tau\,S_{ik}\bigr{)}},\;\;\tau=\frac{1}{\mathbf{T}},\;\mathbf{T}>0.$ (6)
From Eq.(6), a higher $\tau$ increases the model’s confidence, whereas a lower $\tau$ yields a smoother distribution, with greater uncertainty. To ensure that only confident predictions propagate downstream, we filter uncertain detections through an energy-based uncertainty quantification score, $\mathcal{E}(\mathcal{M}_{i})=-\text{log}\sum_{j}\exp\!\bigl{(}S_{ij}\bigr{)}>\tau_{e}$ by rejecting masks exceeding the defined energy threshold $\tau_{e}$. In reality, the environment’s perception quality depends heavily on the accuracy of the observation sources. To account for potential degradations of the robot’s task environment (e.g., low light, occlusions, motion blur), we introduce a function $\Theta_{ij}(\Omega)$ that incorporates environmental conditions for degradation-aware weighting. We downweight probabilities for masks in degraded regions using $\Theta_{ij}(\Omega)=\exp\bigl{(}-\beta\,\|\Omega_{i}\|\bigr{)}$, where $\|\Omega_{i}\|$ quantifies local adverse conditions, and $\beta$ regulates the sensitivity. Formally, Eq. (6) can be rewritten as depicted in Eq. (7):
$p^{\prime}(d_{j}\mid\mathcal{M}_{i},\;\Omega)=\frac{p(d_{j}\mid\mathcal{M}_{i})\cdot\Theta_{ij}(\Omega)}{\sum_{k=1}^{M}\Bigl{(}p(d_{k}\mid\mathcal{M}_{i})\cdot\Theta_{ik}(\Omega)\Bigr{)}}.$ (7)
To spatially ground and track the detected objects, we used the depth map $\mathcal{D}_{t}$. First, at the mask’s centroid $(u_{c}\,,v_{c})$, we compute the depth $z_{c}$ as the median of valid sensor measurements within the local neighbourhood (see Eq. (8)):
$z_{c}=\begin{cases}\mathrm{med}(\mathcal{N}_{r}(u_{c},v_{c})\odot\mathcal{D}_{t}),&\text{if valid}\\ \mathrm{med}(\mathcal{M}_{i}\odot\hat{\mathcal{D}}_{\text{mono}}),&\text{otherwise}\end{cases},$ (8)
where $\hat{\mathcal{D}}_{\text{mono}}$ is the MiDaS [7] monocular depth prediction. Therefore, for the detected object $o_{j}$ with mask centroid $(u_{c}\,,v_{c})$ at depth $z_{c}$, we apply a pinhole camera model to back-project the pixel into 3D space, $\mathbf{x}_{o}\in\mathbb{R}^{3}$, i.e., $\mathbf{x}_{o}=\Pi^{-1}(u_{c}\,,v_{c},z_{c}).$ We then transform $\mathbf{x}_{o}$ to the robot’s base frame using iterative TF lookups to handle temporal synchronisation. Simultaneously, we used the Kalman filter to track the object poses, modelling the state dynamics as $\mathcal{X}_{t+1}=\mathrm{F}\mathcal{X}_{t}+\mathrm{w}$ to smooth pose estimates and account for motion uncertainty, where $\mathrm{F}$ is the motion model.
Finally, given a linguistic command $c$, we reduce the object navigation to selecting the target object $o^{*}$ that maximises:
$o^{*}=\arg\max_{j}\;p^{\prime}\bigl{(}d_{j}\mid\mathcal{M}_{i}\,,\;\Omega\bigr{)}\cdot\text{sim}(d_{j}\,,c),$ (9)
where $\text{sim}(\cdot)=\text{cos}(f_{\text{text}}(d_{j})\,,f_{\text{text}}(c)\bigr{)}$ is the semantic alignment between $d_{j}$ and the command embedding. The output of Eq.  (9) directs the robot to the Kalman-filtered pose, mapping linguistic references (e.g., “navigate to the detected chair”) to explicit 3D coordinates in the robot’s reference frame.
III-E Action Execution and Control
We translate the high-level intents obtained from the DnPP (Subsection III-C) into physical robot actions through the action execution mechanism (AEM), see Fig. 2, top right. Generally, the AEM manages all the navigation tasks and coordinates the rest of the fundamental robot manoeuvres, sensor-based commands, and safety measures. When the user command requires navigation to coordinates $(x_{g},y_{g},z_{g})$, e.g., “go to the locations (6, 0, 0) and (2, 2, 0)” or to pre-defined goal destinations, e.g., “go to the Secretary’s office”, we rely on a higher-level motion planning stack to accomplish these tasks. First, we employ a highly efficient Rao-Blackwellized particle filter-based algorithm [18] to learn occupancy grid maps from the robot’s operational environment. For details on these probabilistic simultaneous localisation and mapping (SLAM) methods, we refer the reader to [90].
To localise the robot within the learned occupancy grid map, we employed the Adaptive Monte Carlo Localisation (AMCL) algorithm [91]. AMCL is a variant of the Monte Carlo Localisation (MCL) [14] which utilises distributed particle filters (a set of weighted hypotheses) to represent the probable state of the robot in the operational environment. With the robot localized in the learned map, zero- and few-shot goal-directed navigation commands become interpretable by the AEM. For navigating the robot to goal coordinates, we employed the MoveBase package of the ROS [72] navigation planner to handle both the path planning and obstacle avoidance, including dynamic and static obstacles.
Beyond the large-scale navigation, we also utilised the AEM to execute low-level movement primitives that do not involve mapping, path planning, or obstacle avoidance. For example, commands like “move in a rectangular pattern of length $3~{}m$ and breadth $2~{}m$ at $0.5~{}m/s$ or “perform a $180^{\circ}$ arc of radius $2~{}m$ are translated into continuous linear and angular velocity profiles. These profiles are then directly mapped into actionable commands for the robot through twist messages, i.e., $\Lambda:(\mathcal{A}_{n}(\phi_{n}),\mathcal{V}_{s})\mapsto\{(\mathbf{v}(t),\mathbf{\omega}(t))\}_{t=1}^{T_{i}}$, where $\mathbf{v}(t)$ and $\mathbf{\omega}(t)$ are the linear and angular velocities respectively, and $T_{i}$ is the time horizon for that action. Additionally, for query-oriented commands that do not involve physical movements (e.g., “send me your coordinates”, “send a photo of your current view” or “describe your surroundings”, etc), the AEM directly queries the robot’s odometry and the visuo-lingual pipeline to capture an image or generate the sensor-based data.
IV Experiments
We conducted experiments in both simulated and real-world environments to validate the full potential of ReLI. In this section, we describe our experimental protocols and present quantitative and qualitative observations gleaned from them.
IV-A Experiment Platforms
We used two robot platforms: (i) a wheeled differential drive robot [63], and (ii) a Unitree Go1 quadruped. Both robots were equipped with an RGB-D camera and a LiDAR sensor for visual and spatial observations. We ran all the simulation experiments with a ground station PC with Nvidia GeForce RTX 3060 Ti GPU, and ROS Noetic distribution. We used Gazebo, an open-source physics-based robotics simulator environment. The environment consists of 11 rooms and an external corridor which replicates a physical indoor office layout, with dimensions $\approx 18.43\times 20.33~{}m$ and $\approx 6.18\times 36.22~{}m$ respectively. We used our PC’s inbuilt microphones for all the experiments involving vocal or audio instructions $c^{v}$.
In the real-world experiments, we used a Lenovo ThinkBook Intel Core i7 with Intel iRIS Graphics. We experimented in our robot station laboratory, measuring $\approx 28.72\times 12.75~{}m$. Both the real-world and simulation environments contained typical real-world furnishings such as tables, chairs, and standing obstacles. We experimented with different LLMs, specifically LLaMA 3.2 [93] and two OpenAI GPT [66] models, namely GPT-4o and GPT-4o-mini. However, among these models, GPT-4o exhibited superior performance in understanding the context of the user’s instructions and the intended actions. Therefore, we utilise it to obtain all the results in Section V.
IV-B Benchmark Design and Datasets
Ultimately, we are mostly interested in the number of languages that ReLI can ground into real-world robotic affordances. For this, we conducted an extensive multilingual evaluation of ReLI to investigate its generalisation across languages. We randomly chose $140$ representative languages from the ISO 639 [27] language catalogue, distributed across the continents. We categorised them based on their resource tiers (i.e., high, low, and vulnerable) and the language family (e.g., Indo-European, Afro-Asiatic, Austro-Asiatic, Sino-Tibetan, Niger-Congo, etc.). Fig. 4 shows the distribution of the language families and their corresponding resource tiers (bottom left).
Similar to the taxonomy in NLLB [89] and Joshi et al. [31], we consider languages with strong digital presence (large-scale corpora, well-established tokeniser, and ISO 639 standards [27]) as high-resource languages (HRL). In contrast, we consider those with a limited digital presence, low-scale training corpora, and less established institutional support as low-resource languages (LRL). Furthermore, we group creoles, vernaculars and rare dialects that have minimal or no recognised status (e.g., susceptible to external pressures, near-extinct or with the UNESCO endangerment status [55, 45]) yet are decodable by LLMs as vulnerable languages (VUL). Figure 4 (top and bottom right) shows the distribution of selected languages across continents, along with approximate representative speakers for the top 15 HRL, LRL, and VUL.
Figure 4: Distributions of the 140 representative languages utilised for ReLI benchmarking. We prioritize the inclusion of low-resource and vulnerable languages in our selection criteria, as we posit that this will rigorously evaluate the robustness and efficacy of our framework (bottom left). Further, to promote inclusive and accessible HRI, we ensured that our selected languages are strategically distributed across the world’s continents (top).
Task instructions and rationales
To ensure a robust benchmark that captures the real-world complexity across the languages, we designed task instructions (see Appendix -C, Table IX) that test ReLI’s multilingual parsing, environment-based decisions, numeric reasoning, conditional branching, etc. Each instruction tests unique combinations of motor primitives, sensor queries, and interactions.
However, while we were unable to quantify all the language-conditioned task instructions that ReLI can perform in real-world, we instead formally structured them at the task level, characterised by the tuple $\mathcal{T}_{T}^{Re}=(\mathcal{G}_{n},\mathcal{W}_{c},\mathcal{Q}_{i},\mathcal{O}_{n},\mathcal{C}_{r})$. Here, $\mathcal{G}_{n}$ includes all the zero-shot spatial or goal-directed navigation tasks (e.g., “navigate to the coordinates $(x_{g},y_{g},z_{g})$ or to a named destination, “head to the kitchen”). $\mathcal{W}_{c}$ are movement commands that involve no direct location targeting, robot localisation or obstacle avoidance (e.g., “move forward $d$ meters at a speed of $v~{}m/s$, “rotate $\theta$ degrees”, etc. $\mathcal{Q}_{i}$ include general information or causal queries (e.g., “which company built you?”, “what are your capabilities?”), and visuo-lingual reasoning tasks (e.g., “what can you see around you?”, “could you send me a photo of your current view?”, etc). $\mathcal{O}_{n}$ includes zero- and few-shot object navigation capabilities (e.g., “go towards the detected chair”). $\mathcal{C}_{r}$ are contextual and descriptive reasoning abilities that allow the robot to interpret commands that require an understanding of context or implicit references. For example, the command “among your known destinations, take me to where I can cook food,” implies navigating to the kitchen, while “take me to where I can handle administrative tasks or inquiries” means head to the secretary’s office.
Figure 5: Distribution of the task instructions utilised for ReLI benchmarking (see Table IX). We refer to task instructions that require the execution of atomic actions as short-horizon tasks. In contrast, those that require strategic planning and the user’s approval or rejection of the generated action plans are considered long-horizon tasks. The labels refer to $G_{n}$ (zero-shot spatial and goal-directed tasks), $W_{c}$ (movement commands with no location targeting), $Q_{i}$ (general information and causal queries), $O_{n}$ (zero- and few-shot object navigation), and $C_{r}$ (contextual and descriptive reasoning abilities).
Fig. 5 shows the distribution of the task instructions. For each language, we conducted $130$ trials (i.e., $130$ random short and long-horizon tasks) covering a balanced mix of the five task-level categories. These resulted in the logged interaction data spanning over $70$K multi-turn conversations. To obtain the task instructions in non-English languages, we utilised the advanced reasoning and multilingual generalisation capabilities of GPT-o1 [12] for interlingual translations. We made this choice to cover languages currently unsupported by Google’s MNMT [29] and NLLB [89] services, e.g., Cherokee, Bislama, African Pidgin, etc. However, to verify the quality and accuracy of our translations, we benchmarked them against the NLLB-200 [89] baseline across 42 languages. We employed multi-dimensional evaluation methods, e.g., BLEU [67], BERTScore [105], etc., to measure both the lexical similarity, semantic fidelity, and safety scores. The comparative results (see Appendix -C, Fig. 12) showed no significant difference (near-equal lexical similarity scores and $>87\%$ in semantic alignments) in the interlingual translations between both models.
Human raters and demongraphics
In addition to the task instructions which we directly provide without a third party, we intermittently recruited $20$ external human-raters (average age of $25\pm 3$, and gender distribution, 70% male, 25% female, and 5% others) fluent in the languages (see Appendix -B, Table VIII) to interact with the robots using either vocal or textual instructions modalities. We instructed the external raters to command the robots to navigate to locations, identify objects, or make general inquiries about the robot’s status and capabilities in their native language. We logged the interaction data which includes the tuple $\mathcal{D}=\{(c_{n},t_{n}^{\mathrm{ins}},t_{n}^{\mathrm{res}},\mathcal{A}_{n},\hat{\mathcal{A}}_{n},s_{n})\}_{n=1}^{N}$, where $c_{n}\in\mathcal{C}_{T}$ is the user’s natural language command, $t_{n}^{\mathrm{ins}}$ is the timestamp when $c_{n}$ was issued, $t_{n}^{\mathrm{res}}$ is the timestamp when the robot began executing $\hat{\mathcal{A}}_{n}$, $\mathcal{A}_{n}$ and $\hat{\mathcal{A}}_{n}$ are the ground-truth and predicted action sequences. $s_{n}$ is a binary action execution success or failure indicator, and $N$ is the total number of task instances.
IV-C Evaluation Metrics
We evaluated ReLI across two dimensions, i.e., quantitative and qualitative. Quantitatively, we assess (i) the accuracy and robustness in multilingual instruction parsing, (ii) the reliability of the action execution mechanism, and (iii) the overall responsiveness and adaptability of the robot’s behaviours. Thus, we defined the following key metrics:
Instruction Parsing Accuracy (IPA)
We define the IPA as the fraction of the user-issued instructions that ReLI correctly parsed to their intended action sequence. Formally, for $N$ commands, we compute $\text{IPA}=\frac{1}{N}\sum_{n=1}^{N}\delta(\mathcal{A}_{n},\hat{\mathcal{A}}_{n})$, where $\mathcal{A}_{n}=\{a_{i}\}_{i=1}^{k}$ is the ground truth action sequence for command $c_{n}$, and $\hat{\mathcal{A}}_{n}=\{\mathcal{A}_{j}(\phi_{j})\}_{j=1}^{n}$ is the parsed action sequence. If $\hat{\mathcal{A}}_{n}$ semantically matches the intended sequence, that is, $\hat{\mathcal{A}}_{n}=\mathcal{A}_{n}$, then $\delta(.)=1$ and 00 otherwise. Notably, we considered partial matches acceptable (e.g. minor parameter discrepancies in speed or distance to the intended goal coordinates) to account for real-world sensor noise.
Task Success Rate (TSR)
We quantify the proportion of trials where the robot completes the intended task within acceptable error thresholds (e.g., within ±$\pm 0.2~{}m$ of navigation to a goal). For a total of $N$ tasks (e.g. navigation to a goal, data request, etc.), we compute: $\text{TSR}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\text{task}}(\hat{\mathcal{A}}_{n},\mathcal{A}_{n})$, where $\delta_{\text{task}}(.)$ indicates success. We considered a task ($n\in\{1,\dots,N\}$) successful if the resulting robot action meets the intended goal (e.g., reaching the specified goal coordinates).
Average Response Time (ART)
We measure the latency from command issuance to the robot’s response with the ART metric. Formally, we compute: $\text{ART}=\frac{1}{N}\sum_{n=1}^{N}(t_{n}^{\text{res}}-t_{n}^{\text{ins}})$, where $t_{n}^{\text{ins}}$ is the time when the instruction is issued and $t_{n}^{\text{res}}$ is the time the robot responds to the instruction.
V Results
In this section, we summarise ReLI’s performance across the key metrics. All experiments were conducted indoors with standard obstacles (tables, chairs, corridors, etc.) and included instructions in the languages described in Section IV-B.
V-A Quantitative Results
[…]
High resources languages (Table I)
In terms of specific language observations, ReLI handled instructions in English, Spanish, and a few others nearly perfectly, with an average IPA $>$ 99%. We attribute this high performance primarily to their large training corpora and well-established linguistic resources, which enhanced the pre-trained LLM prediction accuracy and action parsing. Conversely, certain languages, e.g., Arabic, Chinese, etc, lagged slightly behind other Indo-European high-resource languages. This discrepancy is attributed to the complexities associated with inputting the logographic characters in our interaction interface. Obviously, other than the vocal or audio commands, we relied completely on the translated task instructions for such languages. Furthermore, the TSR for English and Spanish remained consistent with its highest IPA. French, German, etc, also remained slightly above 97% accuracy. Most Indo-European languages, along with Chinese, Japanese, Swahili, Malay, etc., maintained average response times between $2.10-2.20$ seconds, which is ideally a rapid response time for a multilingual system.
TABLE I: Benchmark performance of ReLI on HRL. Accuracies are averaged, std. dev. are within ±$\pm 0.1$. See Appendix -A for details.
Family Indo-European Sin-Ti Afr-As Japo Nig-Co Austr Turk
Lang. English Spanish French German Hindi Russian Portug. Chinese Arabic Japanese Swahili Malay Turkish
IPA (%) 99.6 99.2 98.8 97.7 93.8 96.2 96.9 93.8 92.3 94.6 93.1 95.4 93.8
TSR (%) 99.5 99.0 98.6 97.5 93.6 96.1 96.8 93.7 92.1 94.4 92.9 95.2 93.7
ART (s) 2.10 2.12 2.13 2.14 2.19 2.15 2.15 2.13 2.27 2.18 2.20 2.17 2.18
Legends: Sin-Ti $\rightarrow$ Sino-Tibetan. Afr-As $\rightarrow$ Afro-Asiatic. Japo $\rightarrow$ Japonic. Nig-Co $\rightarrow$ Niger-Congo. Austr $\rightarrow$ Austronesian. Turk $\rightarrow$ Turkic.
Low resource languages (Table II)
ReLI achieved near-high resource performance for IPA and TSR in most LRL, e.g., Irish, Sicilian, Shona, Yoruba and Javanese, all $>$ 96%. However, others, e.g., Serbian, Faroese, Hausa, Amharic, Fijian, Lao, and Quechua are comparatively lower with IPA and TSR $<$ 95%. The ART 2.122.76absent2.122.76$\approx$ 2.12–2.76≈ 2.12 – 2.76s is not drastically higher than the HRL counterparts. Nonetheless, ReLI maintained a reasonably high accuracy and success rate (92 – 98%) in the majority of the low-resource languages.
TABLE II: Benchmark performance of ReLI on LRL. Accuracies are averaged, std. dev. are within ±$\pm 0.1$. See Appendix -A for details.
Family Indo-European Afro-Asiatic Niger-Congo Austronesian Kra-Dai Quechua
Lang. Irish Serbian Faroese Sicilian Hausa Amharic Shona Igbo Yoruba Fijian Javanese Lao Quechua
IPA (%) 97.7 87.7 94.6 96.5 91.5 93.1 96.9 95.4 96.2 90.8 96.9 93.9 92.3
TSR (%) 97.5 87.7 94.5 96.3 91.4 93.0 96.8 95.3 96.0 90.6 96.9 93.7 92.1
ART (s) 2.17 2.76 2.49 2.20 2.23 2.31 2.22 2.24 2.17 2.29 2.12 2.32 2.22
Vulnerable languages (Table III)
ReLI remained robust, even for creoles and vernaculars that typically have fewer or virtually no computational resources and recognised status. It maintained an average IPA and TSR above 94%. This shows the ReLI’s strong capacity to parse and execute instructions in languages with limited digital resources. For instance, Nigerian Pidgin, Tok Pisin, and Haitian Creole approached near-HRL performance, which indicates the ReLI’s ability to utilise their lexical overlap with English or French.
In contrast, some Creoles, e.g., Bislama, exhibited slightly lower IPA and TSR scores, due to their smaller or less standardised corpora. Moreover, Breton, Tiv, and Cherokee highlight the challenges inherent in truly limited resources. Both showed somewhat lower IPA/TSR alongside higher response times (e.g., ART $>2.4$s). Nonetheless, the overall performance across these languages remained highly impressive, showing ReLI’s capacity to handle diverse linguistic typologies despite limited resources.
TABLE III: Benchmark performance of ReLI on vulnerable languages. Accuracies are averaged, std. dev. are within ±$\pm 0.1$. See Appendix -A.
Family Creoles Indo-European Nig-Co Iroq Austr Hm-Mi Turk
Lang. Nig. Pidg. Tok Pisin Bislama Haitian Ossetian Breton Cornish Tiv Cherokee Chuukese Hmong Chuvash
IPA (%) 98.1 95.0 91.9 96.2 94.2 92.3 95.4 91.5 93.1 95.8 97.7 95.4
TSR (%) 97.9 94.8 91.7 96.1 94.0 92.1 95.2 91.3 92.9 95.7 97.6 95.2
ART (s) 2.14 2.21 2.38 2.33 2.23 2.49 2.71 2.67 2.53 2.26 2.28 2.23
Legends: Nig. Pidg. $\rightarrow$ Nigerian Pidgin. Nig-Co $\rightarrow$ Niger-Congo. Iroq $\rightarrow$ Iroquoian. Austr $\rightarrow$ Austronesian. Hm-Mi $\rightarrow$ Hmong-Mien. Turk $\rightarrow$ Turkic.
Impact of instruction horizons on ReLI
We investigate whether short- and long-horizon instructions significantly impact ReLI’s instruction parsing and action execution capabilities. For this, we tested ReLI’s action execution success rate based on the individual task instructions.
Fig. 6 shows the results across selected languages.
Figure 6: TSR across languages and task instructions (top), along with short- and long-horizon performance comparison (bottom). ReLI maintained robust, language-agnostic execution accuracy near and above 90–95% for most tasks.
Notably, as shown in Fig. 6 (top), ReLI achieved nearly 100% success on task instructions involving contextual and descriptive reasoning abilities ($C_{r}$). Causal queries and sensor-based information retrieval tasks ($Q_{i}$) also achieved above 90% success rate in all the tasks. The remainder errors stemmed from instruction ambiguities, especially with insufficient context, which occasionally led to misinterpretation of the user’s intent.
For the goal-directed navigation tasks ($G_{n}$), ReLI achieved above 86% success, with the minority failures due to the navigation planner and partial SLAM errors. The low performance in the object navigation tasks ($O_{n}$) is mostly due to some ambiguous task instructions, which often cause misidentification and navigation to objects based on their descriptions, especially when similar objects exist.
In terms of task horizons, short-horizon tasks (Fig. 6 (bottom right)) exceeded 90% success, compared to the long-horizon counterparts (bottom left). This is consistent with the expectations that pre-trained LLMs easily interpret single-step instructions. Overall, ReLI maintained a high degree of task execution success for both task horizons.
V-B Qualitative Results
While the quantitative evaluation (Section V-A) showed impressive results, it does not fully capture the quality of ReLI’s behaviour. To this end, we collected subjective feedback from the human raters through a 5-point Likert scale survey (1 = strongly unfavourable, 5 = strongly favourable). We gathered the raters’ anecdotal perspectives from a verbal assessment of ReLI’s performance. Specifically, we assessed (i) responsiveness, i.e., perceived latency and promptness of our framework, (ii) correctness and naturalness, and (iii) language-induced performance gap.
Figure 7: Notable human raters’ feedback on ReLI. The majority of the raters assigned favourable (4 – 5) scores for the ease of interaction, comfort/naturalness (V. Comf. $\rightarrow$ Very Comfortable, Uncomf. $\rightarrow$ Uncomfortable), and responsiveness. More than 85% indicated that they could not observe any language-induced gap in the ReLI’s performance.
Fig. 7 shows the notable open‐ended qualitative feedback and the corresponding quantitative ratings from the human raters. With 4 and 5 ratings taken as the most favourable benchmarks, 75% expressed comfort with our framework’s naturalness, and over 85% reported high satisfaction with the robot’s responsiveness to their natural language commands. Further, excluding those that maintained neutrality, none perceived any language‐induced gap that interfered with their instruction execution. Overall, the raters described the interaction as “intuitive,” “cool,” and “natural.” However, some recommend extending support for advanced behaviours, e.g., performing a specialised dance action (e.g., a quadruped robot), given verbal or textual descriptions of the dance style.
For further details on the raters’ demographics, contributed task instructions, and the real-world visual examples of ReLI’s parsed instructions alongside their corresponding actions in diverse languages, we refer the reader to Appendix -B.
VI Limitations
Although ReLI demonstrates robust performance across diverse languages, it is not without limitations. We acknowledge that ReLI relies on large-scale pre-trained LLMs [66, 93, 13, 88] and multimodal VLMs [73, 46] as the backbone. Consequently, its performance is highly influenced by the robustness of these models (in other words, it inherits their limitations). Due to these models’ autoregressive and stochastic nature, they can occasionally produce inconsistent or hallucinated action sequences [23, 70]. This can result in stochastic behaviour from the robot, particularly in the atomic actions that do not require the user’s approval or rejection prior to execution.
Second, while we were unable to quantify all the languages that ReLI can generalise, languages that are not decodable by LLMs can potentially flaw ReLI’s performance. Such languages could cause ReLI to produce erroneous or misinterpreted action sequences. Testing whether chat fine-tuned LLMs, e.g., ChatGPT, can decode the language would be one way to deal with this. Otherwise, ReLI will struggle to ground instructions within the language context.
Furthermore, for vocal or audio-based commands, ReLI relies on accurate language detection and speech recognition. Code-mixed vocal commands and background noise can degrade both the language detection and the instruction transcription. Although we introduced fallback and manual language selection strategies to mitigate these issues, real-world usage might still experience a drop in success rate for consistently noisy or code-mixed environments. Overcoming these acoustic and random noise challenges requires a deeper integration of adaptive noise-cancellation and accent-robust [34, 71, 57] ASR models. Thus, we reserve these for our future work.
Finally, most LLMs are predominantly served via cloud resources, which introduces latency and network connection-dependence issues. In highly dynamic robot tasks or fast-paced operational domains, e.g., search and rescue, time delays caused by network interruptions or high-volume traffic can degrade ReLI’s responsiveness. Thus, we recommend strong internet connectivity while using our framework.
VII Conclusion
In this work, we introduced ReLI, a multilingual, robot-instructable framework that enables free-form human instructions to be grounded in real-world robotic affordances. Our experiments show that ReLI not only interprets commands in high-resource languages at near-human levels, but also extends its capabilities to low-resource, creoles, and endangered languages. Moreover, we observed reliable performance on both short-horizon and long-horizon tasks. ReLI consistently achieved above 90% success in parsing and executing commands. This shows the potential of our framework to enhance the intuitiveness and naturalness of human-robot interaction in linguistically heterogeneous environments. However, despite these advances, further improvements are possible. Better resilience against ambiguous inputs, reduced cloud dependency, and adaptive noise-cancellation remain open challenges, which we aim to tackle in our future work. We believe ReLI advances inclusive, accessible and cross-lingual human-robot interaction to benefit the diverse global communities.